I am having trouble with the steps to solve the following equation: m-3+2=|-3-8|+7. Thank you for your help!
Answer by nyc_function(2733)

m - 3 + 2 =|-3-8| + 7 =
We are solving for m.

m - 1 = |-11| + 7

m - 1 = 11 + 7

m = 18 + 1

m = 19